OpenStudy (anonymous):

7 ft 8 1/4 inches - 6 ft 8 3/4 inches. I need it to be worked out to please ! Thankyouu(:

OpenStudy (anonymous):

Subtract it.

OpenStudy (anonymous):

i know that thanks. I need help still though.

OpenStudy (anonymous):

do 7ft-6ft than take the 1ft 8 1/4 inch and subtract 8 3/4 inch

OpenStudy (anonymous):

thanks!(:

OpenStudy (anonymous):

OMG. im still confused . Now I have 1 ft 8 1/4 - 8 3/4 ? how do i subtract 8 1/4 from 8 3/4 if 3/4 is bigger than 1/4 ?!?!?! LORD HELP ME .

OpenStudy (anonymous):

u need to use the 1ft and turn it to 12 inches and have 20 1/4 inch -8 3/4 inch

OpenStudy (anonymous):

OMG. im still confused . Now I have 1 ft 8 1/4 - 8 3/4 ? how do i subtract 8 1/4 from 8 3/4 if 3/4 is bigger than 1/4 ?!?!?! LORD HELP ME .

OpenStudy (anonymous):

omigod thank you !

OpenStudy (anonymous):

OMG. im still confused . Now I have 1 ft 8 1/4 - 8 3/4 ? how do i subtract 8 1/4 from 8 3/4 if 3/4 is bigger than 1/4 ?!?!?! LORD HELP ME .

OpenStudy (anonymous):

how do i subtract 1/4 from 3/4 thats my main problem .

OpenStudy (anonymous):

if u have 20 1/4 and 8 and 3/4 u can make it 19 5/4 and 8 3/4

OpenStudy (anonymous):

OMG. im still confused . Now I have 1 ft 8 1/4 - 8 3/4 ? how do i subtract 8 1/4 from 8 3/4 if 3/4 is bigger than 1/4 ?!?!?! LORD HELP ME .

OpenStudy (anonymous):

dude, i love you . (: haah thank you so so so so much !

OpenStudy (anonymous):

ya everyone does

OpenStudy (anonymous):

OMG. im still confused . Now I have 1 ft 8 1/4 - 8 3/4 ? how do i subtract 8 1/4 from 8 3/4 if 3/4 is bigger than 1/4 ?!?!?! LORD HELP ME .

OpenStudy (anonymous):

i have like 2 more questions for you ! please ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

OMG. im still confused . Now I have 1 ft 8 1/4 - 8 3/4 ? how do i subtract 8 1/4 from 8 3/4 if 3/4 is bigger than 1/4 ?!?!?! LORD HELP ME .

OpenStudy (anonymous):

class starts at 10:29am and ends at 11:14am. If you cam in 35 minutes late , what percent of class did you miss ?

OpenStudy (anonymous):

so u need to figure out how long class is

OpenStudy (anonymous):

i got 77.7 % because class is 45 min long and you were only there 10 min so u set up the proportion 10/45=x/100 which is 1000=45x

OpenStudy (anonymous):

45 minutes long now what ?

OpenStudy (anonymous):

look at my math set up the proportion than solve

OpenStudy (anonymous):

wait i messed up u want how long u were not there so it is 35/45 but the answer is the same

OpenStudy (anonymous):

where did u get 1000

OpenStudy (anonymous):

that was 10x100 but u should do 35x100 that was a mistake but the answer is the same

OpenStudy (anonymous):

3500?

OpenStudy (anonymous):

so